# Towards Universal Soccer Video Understanding

### v2

@@ Abstract
As a globally celebrated sport, soccer has attracted widespread interest from fans all over the world. This paper aims to develop a comprehensive multi-modal framework for soccer video understanding. Specifically, we make the following contributions in this paper: (i) we introduce SoccerReplay-1988, the largest multi-modal soccer dataset to date, featuring videos and detailed annotations from 1,988 complete matches, with an automated annotation pipeline; (ii) we present the first visual-language foundation model in the soccer domain, MatchVision, which leverages spatiotemporal information across soccer videos and excels in various downstream tasks; (iii) we conduct extensive experiments and ablation studies on event classification, commentary generation, and multi-view foul recognition. MatchVision demonstrates state-of-the-art performance on all of them, substantially outperforming existing models, which highlights the superiority of our proposed data and model. We believe that this work will offer a standard paradigm for sports understanding research.
@@ 1 Introduction
Figure 1: Overview. We present the largest soccer dataset to date, termed SoccerReplay-1988, and the first vision-language foundation model for soccer, MatchVision, capable of various tasks such as event classification and commentary generation.
@@
The existing literature on soccer video analysis predominantly centers on the SoccerNet series datasets [14, 9, 7], which collectively comprise 500 full-match videos for benchmarking a variety of tasks, including event classification [14, 9], commentary generation [34, 36, 40], etc. Despite this extensive coverage, the focus has primarily been on designing specialized models, tailored to narrow, specific tasks, leading to a significant gap in compatibility among models. Such fragmentation underscores the need for a unified analytical framework that can seamlessly integrate the diverse demands of soccer video understanding, promoting more holistic and scalable solutions.
@@
In this paper, we introduce SoccerReplay-1988, the largest and most comprehensive multi-modal soccer video dataset to date, featuring 1,988 complete match videos with rich annotations, such as event labels and textual commentaries. This dataset serves as a solid foundation for developing advanced models in the soccer domain and establishes a challenging new benchmark for evaluating soccer understanding models. In addition, we have harmonized existing datasets to be compatible with ours, further expanding the data resources available to the field.
@@
Leveraging this dataset, we develop the first visual-language foundation model tailored for various soccer understanding tasks, termed MatchVision. It employs the cutting-edge visual-language foundation model as the backbone, e.g., SigLIP [57]. We further extend the framewise visual features to be spatiotemporal with temporal attentions [3], by training on diverse visual-language tasks on SoccerReplay-1988, as depicted in Figure 1. As a result, MatchVision demonstrates strong adaptability across various tasks, including event classification and commentary generation, serving as a universal and unified framework for comprehensive soccer video understanding.
@@
To summarize, we make the following contributions in this paper: (i) we construct SoccerReplay-1988, the largest and most diverse soccer video dataset to date, comprising videos of 1,988 soccer matches with rich annotations, supported by an automated curation pipeline. This provides a solid data foundation for developing robust and comprehensive soccer understanding models; (ii) we present the first visual-language foundation model for soccer understanding, termed MatchVision, which effectively leverages the spatiotemporal information in soccer videos, and can adapt to various tasks such as event classification and commentary generation, serving as a unified framework for soccer understanding; (iii) we establish more comprehensive and challenging benchmarks based on our dataset, enabling more professional evaluation of soccer understanding models; (iv) extensive experiments and ablation studies have demonstrated the superiority of our data and model across various downstream tasks, achieving state-of-the-art performance on both existing benchmarks and our newly established ones. We believe our work potentially offers a viable paradigm for sports video understanding in the future.
@@ 2 Related Works
Sports Understanding [44] is an evolving field that encompasses multiple research topics and integrates diverse data modalities, covering various tasks such as action recognition [14, 9, 15], commentary generation [34, 56, 51, 37, 40], athlete analysis [41, 54], tactical planning [50], sports health [39], and intelligent refereeing [21, 22]. Additionally, with the rapid development of multimodal large language models (MLLMs), recent efforts [26, 52, 53] have attempted to build more generalized frameworks to uniformly handle a variety of sports understanding tasks.
Visual-Language Models [38, 57, 27, 28, 1] have exhibited remarkable performance across extensive applications like classification, segmentation, image-text retrieval, and image captioning. Furthermore, efforts have also ventured into more challenging video understanding [30, 29, 58] tasks, such as temporal alignment [31, 16], dense captioning [60, 5, 55], and audio description [17, 18, 19]. However, these efforts typically focus on general scenarios, limiting their adaptability to specific professional fields. Thus, to advance further into specialized fields such as soccer, this paper presents the first visual-language foundation model tailored for comprehensive soccer understanding.
Soccer Game Analysis has primarily focused on tasks such as action spotting [14, 34], replay grounding [21, 59], commentary generation [34, 36, 40], player tracking [8, 47], state reconstruction [42], camera calibration [6, 9, 8] and foul recognition [21, 22], as facilitated by the SoccerNet [14, 9, 7, 13] series of datasets, with 500 full-match videos from 2015 to 2017. Unlike existing approaches that target designing specific models for distinct tasks, this paper aims to design a unified multi-modal framework that leverages the spatiotemporal information within videos, to serve as a foundation model for soccer video understanding.
@@ 3 SoccerReplay-1988 Dataset
To establish a solid foundation for soccer understanding, we have constructed the largest soccer dataset to date, termed as SoccerReplay-1988. Here, we first outline our data collection details and provide an overview of the dataset in Section 3.1; followed by elaborating on our automated data curation pipeline in Section 3.2; lastly, in Section 3.3, we present the data statistics and discussion.
@@ 3.1 Dataset Collection
To construct the SoccerReplay-1988 dataset, we retrieve and collect untrimmed, full-match videos from the Internet, encompassing a total of 1,988 matches from six European major soccer leagues and championships [^1] Premier (England), Laliga (Spain), Bundesliga (Germany), Serie-a (Italy), League-1 (France) and UEFA Champions League., spanning the 2014-15 to 2023-24 seasons. For each match, we acquire textual commentaries with second-level timestamps from a sports text live website222flashscore.com, with part of them annotated with specific event types such as corner and goal. Additionally, we also incorporate extensive metadata about the matches, including detailed background information about the games, players, coaches, referees, and teams, providing a solid foundation for future soccer understanding research.
We divide the SoccerReplay-1988 dataset into train, validation, and test sets, containing 1,488, 250, and 250 full-match videos with diverse and comprehensive annotations, respectively. These provide diverse training data for downstream tasks, such as event classification and commentary generation, and establish comprehensive and challenging benchmarks for soccer understanding, which will be further discussed in subsequent sections.
@@ 3.2 Automated Data Curation
Given the potential noise in raw data, such as irrelevant video content, inaccurate timestamps, and incomplete event annotations, we design an automated data curation pipeline for curation, comprising (i) temporal alignment, (ii) event summarization, and (iii) anonymization, as in Figure 2.
Temporal Alignment. Here, we divide the match videos into two halves, each starting at the kick-off moment, and adopt the publicly available temporal alignment model from MatchTime [40], to synchronize textual commentary timestamps with those of video frames.
Event Summarization. For samples without event annotations, we leverage LLaMA-3-70B [11] to summarize the events based on textual commentaries. Concretely, we have expanded the event categories from 17 in SoccerNet [9] to 24 types, for finer-grained soccer understanding, for example, categorizing penalties into scored and missed, and integrating modern soccer regulations like VAR. The resulting 24 event labels include: ‘corner’, ‘goal’, ‘injury’, ‘own goal’, ‘penalty’, ‘penalty missed’, ‘red card’, ‘second yellow card’, ‘substitution’, ‘start of game (half)’, ‘end of game (half)’, ‘yellow card’, ‘throw in’, ‘free kick’, ‘saved by goal-keeper’, ‘shot off target’, ‘clearance’, ‘lead to corner’, ‘off-side’, ‘var’, ‘foul (no card)’, ‘statistics and summary’, ‘ball possession’, and ‘ball out of play’. More details on the prompts can be found in Section B.1 of the Appendix.
Anonymization. Similar to [34], we leverage all person and team entity names from the metadata of SoccerReplay-1988, and replace any matched names in textual commentaries with standardized placeholders, such as “[PLAYER]”, “[TEAM]”, “[COACH]”, and “[REFEREE]”, to ensure consistency across the tasks.
Moreover, our proposed data curation pipeline can seamlessly extend to existing datasets, converting the SoccerNet series [9, 34] into our unified data format, termed as SoccerNet-pro. This expansion further enlarges the standardized datasets available for soccer understanding tasks.
@@ 3.3 Statistics & Discussion
Dataset Statistics. As presented in Table 1, our proposed dataset encompasses 3,323 hours of footage from 1,988 soccer matches, with an average duration of 100.3 minutes per match. The videos range in resolution from 360p to 720p and frame rates between 25 and 30 FPS.
For textual annotations, this dataset features approximately 150K precisely temporal-aligned commentaries, averaging 76 per match. These commentaries cover 4,467 unique words, significantly surpassing the 2,873 words in existing datasets [34, 40], greatly enriching the textual diversity. Automated event summarization based on these commentaries has yielded about 150K event annotations.
SoccerReplay-test Benchmark. To facilitate a more comprehensive evaluation of soccer understanding models, we integrate 250 matches from SoccerReplay-1988 with 50 matches from the curated SoccerNet-pro, establishing a more challenging benchmark, named SoccerReplay-test, for event classification and commentary generation. This benchmark features nearly four times larger than existing datasets and comprises finer-grained event labels, richer textual commentaries, and up-to-date soccer regulations.
Discussion. To summarize, our proposed SoccerReplay-1988 dataset demonstrates advancements in three aspects: (i) it is the largest soccer video dataset to date, with nearly four times more videos than existing datasets; (ii) it features more professional and diverse annotations, more suitable for fine-grained and comprehensive soccer understanding tasks; (iii) it employs an automated curation pipeline for annotations, thus is scalable to make a solid data foundation for future research in soccer understanding.
Figure 3: Overview of MatchVision: (a) The model architecture and its spatiotemporal feature extraction process; (b) Details of visual encoder pretraining, such as supervised training and video-language contrastive learning; (c) Implementation details of specific heads for various downstream tasks, including commentary generation, foul recognition, and event classification.
@@ 4 Method
In this paper, we aim to develop a versatile visual-language foundation model, MatchVision, specifically tailored for analyzing diverse soccer video tasks. We start by outlining our problem formulation in Section 4.1. Next, in Section 4.2, we detail the architecture of MatchVision. The training procedures are thoroughly discussed in Section 4.3. Finally, we describe the configurations for our downstream tasks in Section 4.4, demonstrating the practical applications and effectiveness of our model.
@@ 4.1 Problem Formulation
Here, $\Psi=\{\Psi_{\mathrm{cls}},\Psi_{\mathrm{Cmt}},\Psi_{\mathrm{Foul}}\}$ denotes the task-specific heads, while $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{F}$ refer to the output event types, textual commentaries, and foul types, respectively. This unified framework not only effectively learns relevant spatiotemporal features, but also facilitates seamless integration across various downstream tasks tailored for comprehensive soccer game understanding.
@@ 4.2 Architecture
Token Embedding. In accordance with the convention in Vision Transformer [10], each frame ($\mathcal{I}_{i}$) from the video segment $\mathcal{V}=\{\mathcal{I}_{1},\mathcal{I}_{2},\cdots,\mathcal{I}_{T}\}$ is divided into $M$ non-overlapping patches of size $P\times P$ that span the entire frame.
@@
Spatiotemporal Attention Block. Taking inspiration from TimeSformer [3], we utilize interleaved temporal and spatial attention to integrate the spatiotemporal information in soccer videos. Concretely, each spatiotemporal attention block comprises a temporal self-attention layer and a spatial self-attention layer, i.e., $\phi_{t}(\cdot)$ and $\phi_{s}(\cdot)$, respectively.
Given the video feature $\mathbf{z}\in\mathbb{R}^{T\times(M+1)\times D}$, we alternate the temporal and spatial attention: temporal attention facilitates interactions among tokens at the same spatial positions across distinct frames, while spatial attention enables interactions among tokens within the same frame. Residual connections are employed in each layer. After passing through a total of $\mathbf{K}$ spatiotemporal attention blocks, the resulting feature ($\mathcal{F}$) captures both intra-frame and inter-frame relationships, i.e., $\mathcal{F}=[\phi_{s}(\phi_{t}(\mathbf{z}))]^{\mathbf{K}}\in\mathbb{R}^{T\times(M+1)\times D}$.
Aggregation Layer. To get the video-level features, we employ an aggregation layer on the frame-wise spatiotemporal features. Specifically, for the $i$-th frame, we utilize spatial self-attention to aggregate information into its [cls] token, denoted as ^$\hat{\mathcal{F}}_{i}^{\mathrm{cls}}=\Phi_{\mathrm{Agg}}(\mathcal{F}_{i})$. We then concatenate the [cls] tokens of all frames, to get the final video feature ($\mathcal{F}_{\mathcal{V}}$), that effectively encapsulates the spatiotemporal characteristics of the soccer video segments, and enables it to be applicable for various downstream soccer understanding tasks. This process can be formulated as:

### v3

@@ Abstract
As a globally celebrated sport, soccer has attracted widespread interest from fans all over the world. This paper aims to develop a comprehensive multi-modal framework for soccer video understanding. Specifically, we make the following contributions in this paper: (i) we introduce SoccerReplay-1988, the largest multi-modal soccer dataset to date, featuring videos and detailed annotations from 1,988 complete matches, with an automated annotation pipeline; (ii) we present an advanced soccer-specific visual encoder, MatchVision, which leverages spatiotemporal information across soccer videos and excels in various downstream tasks; (iii) we conduct extensive experiments and ablation studies on event classification, commentary generation, and multi-view foul recognition. MatchVision demonstrates state-of-the-art performance on all of them, substantially outperforming existing models, which highlights the superiority of our proposed data and model. We believe that this work will offer a standard paradigm for sports understanding research.
@@ 1 Introduction
Figure 1: Overview. We present SoccerReplay-1988, the largest soccer dataset to date, and a powerful soccer-specific visual encoder, MatchVision, capable of excelling in various tasks such as event classification and commentary generation.
@@
Existing research in soccer video analysis primarily revolves around the SoccerNet series datasets [15, 10, 7], which comprise 500 full-match videos for benchmarking various tasks, such as action spotting [15, 10] and commentary generation [35, 37, 41]. Despite this extensive coverage, the focus has predominantly been on designing specialized models for task-specific applications, leading to fragmented and incompatible solutions. Such fragmentation underscores the need for a unified framework capable of integrating diverse demands, enabling more holistic and scalable advancements in soccer video understanding.
@@
In this paper, we introduce SoccerReplay-1988, the largest and most comprehensive multi-modal soccer video dataset to date, featuring 1,988 complete match videos with rich annotations, such as event labels and textual commentaries. This dataset offers a solid foundation for developing advanced soccer understanding models and establishes a challenging new benchmark for the field. Additionally, we have harmonized existing datasets to be compatible with ours, further expanding the available data resources.
@@
Leveraging this dataset, we develop MatchVision, an advanced soccer-specific visual encoder tailored for diverse soccer understanding tasks. It employs the cutting-edge visual-language foundation model as the backbone, e.g., SigLIP [60]. We further extend framewise visual features into spatiotemporal representations with temporal attentions [3], by training on diverse visual-language tasks on SoccerReplay-1988, as depicted in Figure 1. As a result, MatchVision exhibits strong adaptability across various tasks, such as event classification and commentary generation, serving as a universal and unified framework for comprehensive soccer video understanding.
@@
To summarize, we make the following contributions in this paper: (i) we construct SoccerReplay-1988, the largest and most diverse soccer video dataset to date, featuring videos of 1,988 soccer matches with rich annotations, supported by an automated curation pipeline. This provides a solid foundation for developing robust and comprehensive soccer understanding models; (ii) we present a powerful soccer-specific visual encoder, termed MatchVision, which effectively leverages spatiotemporal information in soccer videos, and can adapt to various tasks such as event classification and commentary generation, serving as a unified framework for soccer understanding; (iii) we establish more comprehensive and challenging benchmarks based on our dataset, enabling more professional evaluation of soccer understanding models; (iv) extensive experiments and ablation studies demonstrate the superiority of our data and model across various downstream tasks, achieving state-of-the-art performance on both existing benchmarks and our newly established ones. We believe this work offers a viable paradigm for future sports video understanding.
@@ 2 Related Works
Sports Understanding [46] is an evolving field that encompasses multiple research topics and integrates diverse data modalities, covering various tasks such as action spotting [15, 10, 16], commentary generation [35, 59, 53, 54, 38, 41], athlete analysis [42, 57], tactical planning [51], sports health [40], and intelligent refereeing [22, 23]. Furthermore, with the rapid development of multimodal large language models (MLLMs), recent efforts [52, 27, 55, 56] have attempted to build more generalized frameworks to uniformly handle a variety of sports understanding tasks.
Visual-Language Models [39, 60, 28, 29, 1] have exhibited remarkable performance across extensive applications like classification, segmentation, image-text retrieval, and image captioning. Recent efforts have ventured into more challenging video understanding [31, 30, 61, 62, 43] tasks, such as temporal alignment [32, 17], dense captioning [64, 5, 58], and audio description [18, 19, 20]. However, these efforts typically focus on general scenarios, limiting their adaptability to specific professional fields. Thus, this paper aims to bridge this gap by advancing visual-language models tailored for comprehensive soccer understanding.
Soccer Game Analysis [9] has primarily focused on tasks such as action spotting [15], replay grounding [10, 63], commentary generation [35, 37, 41], player tracking [8], state reconstruction [44], camera calibration [6, 10, 8] and foul recognition [22, 23], as facilitated by the SoccerNet [15, 10, 7, 14] series datasets, with 500 full-match videos from 2015 to 2017. Unlike existing methods that target designing specific models for distinct tasks, this paper aims to design a unified multi-modal framework that leverages spatiotemporal information within videos, serving as a specialized visual encoder for soccer video understanding.
@@ 3 SoccerReplay-1988 Dataset
To establish a solid foundation for soccer understanding, we construct SoccerReplay-1988, the largest soccer dataset to date. Here, we first outline our data collection details and an overview of the dataset in Sec. 3.1; followed by elaborating on our automated data curation pipeline in Sec. 3.2; lastly, in Sec. 3.3, we present the data statistics and discussion.
@@ 3.1 Dataset Collection
To construct the SoccerReplay-1988 dataset, we have collected untrimmed, full-match videos from the Internet, encompassing a total of 1,988 matches from six European major soccer leagues and championships [^1] Premier (England), Laliga (Spain), Bundesliga (Germany), Serie-a (Italy), League-1 (France) and UEFA Champions League., spanning the 2014-15 to 2023-24 seasons. For each match, we acquire textual commentaries with second-level timestamps from a sports text live website222flashscore.com, with part of them annotated with specific event types such as corner and goal. Additionally, we also incorporate extensive metadata, including detailed background information about the games, players, coaches, referees, and teams, providing a solid foundation for future soccer understanding research.
We partition the SoccerReplay-1988 dataset into train, validation, and test sets, containing 1,488, 250, and 250 full-match videos with diverse and comprehensive annotations, respectively. These sets provide rich training data for downstream tasks, such as event classification and commentary generation, while establishing comprehensive and challenging benchmarks for soccer understanding, as further discussed in subsequent sections.
@@ 3.2 Automated Data Curation
Given the potential noise in raw data, such as irrelevant video content, inaccurate timestamps, and incomplete event annotations, we design an automated data curation pipeline, comprising (i) temporal alignment, (ii) event summarization, and (iii) anonymization, as illustrated in Figure 2.
Temporal Alignment. Here, we divide match videos into two halves, each starting at kick-off, and adopt the temporal alignment model from MatchTime [41], to synchronize textual commentary timestamps with those of video frames.
Event Summarization. For samples without event annotations, we leverage LLaMA-3-70B [12] to summarize the events based on textual commentaries. Concretely, we have expanded the event categories from 17 in SoccerNet [10] to 24 types, for finer-grained soccer understanding, for example, categorizing penalties into scored and missed, and integrating modern soccer regulations like VAR. The resulting 24 event labels include: ‘corner’, ‘goal’, ‘injury’, ‘own goal’, ‘penalty’, ‘penalty missed’, ‘red card’, ‘second yellow card’, ‘substitution’, ‘start of game (half)’, ‘end of game (half)’, ‘yellow card’, ‘throw in’, ‘free kick’, ‘saved by goal-keeper’, ‘shot off target’, ‘clearance’, ‘lead to corner’, ‘off-side’, ‘var’, ‘foul (no card)’, ‘statistics and summary’, ‘ball possession’, and ‘ball out of play’. More details on the used prompts are provided in the Appendix.
Anonymization. Similar to [35], we extract all person and team entity names from the metadata of SoccerReplay-1988, and replace them in textual commentaries with placeholders, such as “[PLAYER]”, “[TEAM]”, “[COACH]”, and “[REFEREE]”, ensuring consistency across tasks.
Moreover, our data curation pipeline can seamlessly extend to existing datasets, converting the SoccerNet series [10, 35] into our unified data format, termed SoccerNet-pro. This expansion further enlarges the standardized datasets available for soccer understanding tasks.
@@ 3.3 Statistics & Discussion
Dataset Statistics. As shown in Table 1, our dataset encompasses 3,323 hours of footage from 1,988 soccer matches, with an average duration of 100.3 minutes per match. The videos range in resolution from 360p to 720p and frame rates between 25 and 30 FPS.
For textual annotations, this dataset features approximately 150K commentaries, averaging 76 per match, precisely temporal-aligned by the robust alignment model from MatchTime [41]. These commentaries cover 4,467 unique words, significantly surpassing the 2,873 words in existing datasets [35, 41], greatly enriching textual diversity. Automated event summarization based on these commentaries has yielded about 150K event annotations. Notably, a random sampling of 2% of the data yields 98% manual verification accuracy, ensuring high-quality automated labeling.
SoccerReplay-test Benchmark. To facilitate a more comprehensive evaluation of soccer understanding models, we integrate 250 matches from SoccerReplay-1988 with 50 matches from the curated SoccerNet-pro, establishing SoccerReplay-test, a more challenging benchmark for event classification and commentary generation. This benchmark features nearly four times larger than existing datasets and comprises finer-grained event labels, richer textual commentaries, and up-to-date soccer regulations.
Figure 3: Overview of MatchVision. (a) The model architecture and its spatiotemporal feature extraction process; (b) Details of visual encoder pretraining, including supervised classification and video-language contrastive learning; (c) Implementation details of specific heads for various downstream tasks, including commentary generation, foul recognition, and event classification.
Discussion. To summarize, SoccerReplay-1988 exhibits advancements in three aspects: (i) it is the largest soccer video dataset to date, with nearly four times more videos than existing datasets; (ii) it features more professional and diverse annotations, more suitable for fine-grained and comprehensive soccer understanding tasks; (iii) It employs an automated curation pipeline for annotations and is thus scalable to provide a solid data foundation for future research. All data from SoccerReplay-1988, including videos and annotations, are open-source for non-commercial use.
@@ 4 Method
In this paper, we aim to develop a soccer-specific visual encoder, MatchVision, tailored for diverse soccer video analysis tasks. We start by outlining our problem formulation in Sec. 4.1. Next, in Sec. 4.2, we detail the architecture of MatchVision. The training procedures are thoroughly discussed in Sec. 4.3. Finally, we describe the configurations for downstream tasks in Sec. 4.4, demonstrating the practical applications and effectiveness of our model.
@@ 4.1 Problem Formulation
Here, $\Psi=\{\Psi_{\mathrm{cls}},\Psi_{\mathrm{Cmt}},\Psi_{\mathrm{Foul}}\}$ represents the task-specific heads, with $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{F}$ denoting the output event types, textual commentaries, and foul types, respectively. This unified framework effectively learns relevant spatiotemporal features, and enables seamless integration across various downstream tasks for comprehensive soccer understanding.
@@ 4.2 Architecture
Token Embedding. In accordance with the convention in Vision Transformer [11], each frame ($\mathcal{I}_{i}$) from the video segment ($\mathcal{V}=\{\mathcal{I}_{1},\mathcal{I}_{2},\cdots,\mathcal{I}_{T}\}$) is divided into $M$ non-overlapping patches of size $P\times P$ that span the entire frame.
@@
Spatiotemporal Attention Block. Similar to TimeSformer [3], we utilize interleaved temporal and spatial attention to integrate spatiotemporal information in soccer videos. Concretely, each spatiotemporal attention block comprises a temporal self-attention layer and a spatial self-attention layer, i.e., $\phi_{t}(\cdot)$ and $\phi_{s}(\cdot)$, respectively.
Given a video feature ($\mathbf{z}\in\mathbb{R}^{T\times(M+1)\times D}$), we alternate temporal and spatial attention: temporal attention facilitates interactions among tokens at the same spatial positions across distinct frames, while spatial attention enables interactions among tokens within the same frame. Residual connections are employed in each layer. After passing through a total of $\mathbf{K}$ spatiotemporal attention blocks, the resulting feature ($\mathcal{F}$) captures both intra-frame and inter-frame relationships, i.e., $\mathcal{F}=[\phi_{s}(\phi_{t}(\mathbf{z}))]^{\mathbf{K}}\in\mathbb{R}^{T\times(M+1)\times D}$.
Aggregation Layer. To obtain video-level features, we employ an aggregation layer on the frame-wise spatiotemporal features. Specifically, for the $i$-th frame, we utilize spatial self-attention to aggregate information into its [cls] token, denoted as ^$\hat{\mathcal{F}}_{i}^{\mathrm{cls}}=\Phi_{\mathrm{Agg}}(\mathcal{F}_{i})$. Concatenating the [cls] tokens of all frames yields the final video feature ($\mathcal{F}_{\mathcal{V}}$), that effectively encapsulates spatiotemporal characteristics of soccer video segments, thus enabling it to be applicable for various downstream soccer understanding tasks. This process can be formulated as: